### v1

@@ 6 Related Work
Other works have explored architectures and objectives that train the backbone to go beyond next-token prediction. This includes non-autoregressive models (Gu et al., 2018), energy-based models Dawid & LeCun (2023), diffusion models (Gong et al., 2023), and variants of Transformers learning to predict multiple tokens at the same go (Qi et al., ; Monea et al., 2023) or injecting “lookahead” data (Du et al., 2023a). Teacherless training — proposed as “parallel speculative sampling (PaSS)” in Monea et al. (2023) — provides an arguably simple such approach that involves a trivial modification to teacher-forcing. Note that while research in parallel decoding too is concerned with predicting multiple future tokens (Stern et al., 2018), the goal is purely inference-time efficiency — which is also the setting under which Monea et al. (2023) propose teacherless training.
One may argue that reinforcement learning-based training (Ranzato et al., 2016; Wu et al., 2016; Bahdanau et al., 2017; Paulus et al., 2018; Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022) is another way to build backbones that go beyond teacher-forcing. However, it is worth noting that the gradients in these techniques boil down to teacher-forcing on the model’s own generated answer. Furthermore, if we desire that the model be able to generate a solution that can plan ahead of time, it is unclear how a model can go from a complete inability to plan (that may assign near-zero probability to the true plan in an exponential space of solutions), to discovering the correct plan simply through preference-based feedback (see (Havrilla et al., 2024) for related empirical evidence).
Another line of work — spanning language (Bengio et al., 2015; Goyal et al., 2016), imitation learning (Ross et al., 2011; Ross & Bagnell, 2010; 2014) and structured prediction (Daumé III et al., 2009; Chang et al., 2015) — has been aimed at addressing the Snowball Failure, under the assumption that the model has otherwise learned an accurate next-step predictor. Broadly, the idea is to train the model on a mixture of the ground truth sequences and the model-generated sequences themselves, as a way to ensure that the test-time and training-time distributions are as similar as possible. These techniques however do not address the failure to learn a good next-step predictor in the first place.
Such limits of end-to-end learning have been echoed in literature on learning with chain-of-thought-type supervision. Recent theoretical works have shown broad classes of tasks (e.g., any function efficiently computed by a Turing machine) where prepending CoT to the end target allows efficiently learning tasks; yet, there are “multi-hop reasoning” tasks that are unlearnable end-to-end (i.e., without intermediate supervision) either due to computational hardness (Wies et al., 2023) or representational limits (Malach, 2023)). Earlier theoretical works Shalev-Shwartz et al. (2017); Shalev-Shwartz & Shashua (2016) have similarly proven negative results for end-to-end learning in similar settings. Similar empirical arguments have been made in neural network literature (Gülçehre & Bengio, 2016; Glasmachers, 2017) and also more recently, in language models on complex reasoning and math problems (Nye et al., 2021; Ling et al., 2017; Cobbe et al., 2021; Piekos et al., 2021; Zelikman et al., 2022; Recchia, 2021; Cobbe et al., 2021).
It is worth noting though that the above lines of work are concerned with chain-of-thought that is present before the end target; in our setup, this supervision is presented only after the end target. Surprisingly, some of our teacherless models manage to utilize even such hindsight chain-of-thought. This success is not fully explained by existing positive results about chain-of-thought supervision, such as Wies et al. (2023); Malach (2023), where supervision is provided before the end target.
Shortcut-learning in language models. A line of work has empirically and theoretically analyzed how Transformer-based language models learn superficial shortcuts to (partially) solve tasks such as learning multiplication (Dziri et al., 2023), automata (Liu et al., 2023), recursion (Young & You, 2023), reading comprehension (Lai et al., 2021) and multiple-choice questions (Ranaldi & Zanzotto, 2023)
However, these shortcuts must not be confused with the Clever Hans cheating induced by teacher-forcing. First, these aforementioned shortcuts exist independent of teacher-forcing: these are correlations between the prefix (such as the initial digits of two multiplicands) and the final answer (the initial digits of the product) in the underlying training distribution. But Clever Hans cheats arise only upon teacher-forcing as they are correlations between the prefixes of the answer itself to the rest of the answer. Second, the above shortcuts only fail out-of-distribution (such as when the number of multiplied digits is increased, where the failure is in length generalization (Anil et al., 2022)). In contrast, the Clever Hans cheat is much more severe as it causes in-distribution failure. Thirdly, the aforementioned empirical observations are specific to Transformers, and the theoretical arguments rely crucially on properties of the Transformer (such as its non-recurrence and convolution, or its self-attention modules). Our argument however only relies on the teacher-forcing objective with no reliance on the Transformer architecture, and is demonstrated even for the recurrent Mamba architecture.
Please see § F for more related works.
7 Limitations
We note that our arguments are empirical and conceptual. We have not provided a formal proof for our arguments. We have also not demonstrated failure for very large models such as Llama2 (Touvron et al., 2023) or Mistral (Jiang et al., 2023). Next, beyond the minimal path-finding setting, we have not demonstrated or characterized the range of problems where teacher-forcing-induced failure may occur. We only intuitively believe it should extend to other problem-solving tasks and creative-writing tasks that require lookahead. It is also unclear if it generalizes to run-of-the-mill text-generation tasks.
@@ 8 Conclusion
Next-token prediction lies at the heart of modern language models which have demonstrated tremendous empirical success in a range of general tasks. Theoretically too, we know by the chain rule of probability that, next-token predictors can express any imaginable distribution over tokens. Thus, it is tempting to view next-token prediction as a formidable approach to modeling language and intelligence. Our work crystallizes the core arguments around why this optimism may be misplaced.
First, we emphasize not to conflate the two modes of next-token prediction: autoregressive inference and teacher-forced training. While existing criticisms primarily challenge autoregressive inference, they assume that teacher-forcing learns a good next-token predictor. We challenge this very assumption, finding that even in a straightforward task, there is failure due to teacher-forcing — not due to autoregressive inference or the architecture. This casts a shadow over more complex tasks. For instance, as we speculate in § B, can a model trained to predict the next token of tens of thousands of fiction novels, learn to generate plot twists?
@@
9 Impact
@@
Acknowledgments: We would like to thank Colin Raffel, Tiago Pimentel, and Surbhi Goel for their extensive feedback on a draft of this preprint, especially for pointers to some key references.
Appendix A Teacher-forcing Failure and Snowballing Failure are distinct
@@
Appendix B An illustration via story-telling
@@
Appendix C More on Teacherless Training
Recall that our hypothesis is that the teacherless model automatically learns to fit the targets in the reverse order, since the path from the ${v}_{\tt{goal}}$ is unique. This is indeed what we find in Fig 4, where the accuracies of the later tokens become higher earlier.
@@ Appendix C More on Teacherless Training
Note though that this is a fairly difficult computation to implement. First, when the model predicts $v_{i}$, it must require the identify of $v_{i+1}$. However, this identity is not fed as input to the model, in the absence of the teacher. Thus the model must have computed $v_{i+1}$ and crucially, stored that in one of its its internal representations. Then, by induction, when predicting the first node $v_{1}$, the model must know the identity of all the other nodes in the path. In other words, the model must have (a) computed and (b) stored the whole solution in its hidden representations before it outputs the first token. This is a substantial type of lookahead that some of our models are able to achieve under teacherless training.
Appendix D More Experimental Results
D.1 Snowball Failure
@@
D.2 Clever Hans Cheating Accuracies
@@
D.3 More Detailed Accuracies
@@
Appendix E Other experimental details
E.1 Tokenization
@@ E.2 Models
Appendix F More Related Work
@@
Other arguments about next-token prediction We survey related arguments of next-token prediction, orthogonal to our main discussion regarding planning. Allen-Zhu & Li (2023); Lv et al. (2023) report that language models that are trained on A equals B are unable to infer B equals A, which Allen-Zhu & Li (2023) suggest is due to autoregressive left-right training. Du et al. (2023b); Welleck et al. (2020) formalize the limitation that autoregressive models may potentially assign non-zero probability to infinite-length strings, thus leading to non-terminating inference. Li et al. (2024) provide a Transformer-specific analysis of how self-attention affects the optimization geometry of next-token prediction. Thrampoulidis (2024) provide an analysis of the implict bias of optimization with next-token prediction for linear models.
@@
Other limitations of Transformers Merrill & Sabharwal (2023a) identify limitations of the representative power of Transformer architecture when the arithmetic precision is logarithmic in the number of input tokens. Bender et al. (2021) criticize GPT-like language models as simply parroting out training data with minor stochasticity, while Arkoudas (2023) report that such models struggle with reasoning, even if not a stochastic parrot. Young & You (2023) study masked language (T5, BERT) models (not causally-trained) and argue there are inconsistencies in the probabilities that they assign. E.g., when conditioned on ‘white’, the probability of ‘rice’ may be higher ‘bread’ but the probability of ‘white bread’ and ‘white rice’ are the opposite. Artetxe et al. (2022) empirically analyze the effect of bidirectional attention and bidirectional supervision (as in masked language modeling) during pretraining on the ability of the model to do various things, including next-token prediction. Springer et al. (2024) argue that autoregressive Transformers compute sub-optimal embeddings that can be improved by repeating the input text twice.
@@
Going beyond next-token prediction-based training. Finally, we note that some works (Gurnee et al., 2023; Meng et al., 2022; Pal et al., 2023) aim to recover future tokens that an already-trained model may predict based on the internal layers of the current token. Note that the success of this does not imply that the model necessarily plans well. This only means that it is possible to recover what the already-trained model wants to generate in the future (which may simply be a bad plan). Pfau et al. (2023) train a language model to predict in reverse with the orthogonal goal of finding prefixes that elicit certain behaviors.

### v2

@@ 6 Related Work
Going beyond next-token prediction. Various works have explored architectures and objectives that train the backbone to go beyond next-token prediction. This includes non-autoregressive models (Gu et al., 2018), energy-based models (Dawid & LeCun, 2023), diffusion models (Gong et al., 2023), and variants of Transformers learning to either predict future tokens in a different ordering (Li et al., 2021) or all at one go (Qi et al., 2020; Monea et al., 2023), or injecting “lookahead” data (Du et al., 2023a). Teacherless training — proposed as “parallel speculative sampling (PaSS)” in Monea et al. (2023) — provides an arguably simple such approach that involves a trivial modification to teacher-forcing. Note that while research in parallel decoding too is concerned with predicting multiple future tokens (Stern et al., 2018), the goal is purely inference-time efficiency — which is also the setting under which Monea et al. (2023) propose teacherless training.
@@ 7 Conclusion
Next-token prediction lies at the heart of modern language models which have empirically demonstrated tremendous success in wide-ranging tasks. Theoretically too, we know by the chain rule of probability that, next-token predictors can express any distribution over tokens. It is tempting then to view next-token prediction as a formidable approach to modeling intelligence. Our work crystallizes the core arguments around why this optimism may be misplaced.
We emphasize not to conflate the two modes of next-token prediction: autoregressive inference and teacher-forced training. While existing criticisms primarily challenge autoregressive inference, they assume that teacher-forcing learns a good next-token predictor. We challenge this very assumption, finding that even in a straightforward task, there is failure due to teacher-forcing — not due to autoregressive inference or the architecture. This casts a shadow over more complex tasks. For instance, as we speculate in §C, can a model trained to predict the next token of thousands of fiction novels, learn to generate plot twists?
@@
We point the reader to §A for a discussion of the limitations of our study.
Acknowledgments
We would like to thank Colin Raffel, Tiago Pimentel, and Surbhi Goel for their extensive feedback on a draft of this preprint, especially for pointers to some key references.
Impact Statement
@@ Acknowledgments
Appendix A Limitations
  1. 1.
    Our arguments are empirical and conceptual. We have not provided a formal proof for our arguments.
  2. 2.
    We have also not demonstrated failure for very large models such as Llama2 (Touvron et al., 2023) or Mistral (Jiang et al., 2023).
  3. 3.
    We note that there may be specific workarounds to make the Transformer (efficiently) learn the path-star task. For example, it may become solvable with other pre-trained models which may have been taught path-finding with step-by-step supervision (see Remark 1). The task may also be solvable via other workarounds such as in-context learning or multi-modal learning (where the model visually processes the image of the graph). Furthermore, since the path-star problem composes the same subroutine over itself, it allows efficient parallelizable solutions Sanford et al. (2024) that may not require learning a sequential composition of discrete subroutines. This can break the assumption of Proposition 3, making the optimization tractable. We warn the reader that these are however only specialized workarounds for this specific task; our broader point is that (a) next-token-prediction/teacher-forcing may still be relatively inefficient compared to teacherless training and (b) there may still be other novel tasks not seen during pre-training, or not solvable visually, or where no paralleizable solutions exist, for which similar failures may occur. We discuss this in detail in Remark 1.
  4. 4.
    Nevertheless, beyond the minimal path-finding setting, we have not demonstrated or characterized the range of problems where teacher-forcing-induced failure may occur. We only intuitively believe it should extend to other problem-solving tasks and creative-writing tasks that require lookahead (see § C).
  5. 5.
    It is also unclear if this failure generalizes to run-of-the-mill text-generation tasks.
Appendix B Teacher-Forcing Failure and Snowballing Failure are Distinct
@@
Appendix C An Illustration via Story-Telling
@@
Appendix D Other Remarks
Remark 1.
(Conditions under which first token becomes decipherable end-to-end) There are certain corner cases where teacher-forcing can learn the (otherwise indecipherable) first token efficiently, without having to brute-force search an exponential space of algorithms. We enumerate these below. (Note though that even if the problem does become tractable, it can still be the case in these problems, next-token prediction is less efficient than predicting multiple future tokens.)
  1. 1.
    Lucky prior biases: If the model happens to have been exposed to certain relevant kinds of supervision during pre-training, then the model will be biased towards a favorable part of the search space, and chance upon the right algorithm much quicker.
    1. (a)
      If the model had witnessed the same task but with the correct step-by-step supervision, then the prior would assign high probability to the correct algorithm. (One can imagine that the the true end-to-end algorithm itself becomes a readily available subroutine in this case.)
    2. (b)
      Or, in the specific path-star example, if the prior bias assigns high probability to all $l$ subroutines being identical, then the search only needs $O(|\mathcal{C}|)$ time (where $\mathcal{C}$ is the set of candidate subroutines.) For illustration, see the Transformer construction for k-hop problems in (Sanford et al., 2024).
    Note that in such a case, one may still be able to demonstrate intractability by constructing slight variations of the tasks that defy such prior biases (e.g., tasks where the subroutines are not identical).
  2. 2.
    Small graph size. If the number of edges is very small (say $|E|$) in proportion to the training data, then the model can learn alternative solutions that “memorize” the problem:
    1. (a)
      Naive memorization: If the vocabulary has only $|\mathcal{V}|$ possible node ID’s, then there are only $O(|\mathcal{V}|^{|E|})$ possible adjacency lists the model can see. So, if the training data has at least $\Omega(|\mathcal{V}|^{|E|})$ datapoints, then, every test example is seen with high probability during training. Here, the model merely needs to look up exact replicas from training, and regurgitate the subsequent values from the training string.
    2. (b)
      Node-ID-agnostic memorization: If the number of training data is $\Omega(|E|!)$, the model can still implement a form of memorization, but this requires a cleverer strategy that is agnostic to the node ID’s. Concretely, for a fixed assignment of node ID’s, there are only $O(|E|!)$ ways to permute the adjacency list. Assume that the model sees all such permutations during training (albeit with varying instantiations of the node IDs). During inference, the model can first look up whether the test input corresponds to an existing permutation it had witnessed (possibly with different node ID’s). For example $1\to 2;2\to 3;4\to 1$ would correspond to $10\to 20;20\to 30;40\to 10$. Then the model merely needs to recall from its memory, the index where the target token is located in the input for that permutation. The model can then look at the same index in the test input and output the node ID located there. Thus, if the target token was $4$ for the first sequence above, the model can output $40$ for the other sequence.
    Note that this doesn’t contradict Proposition 3 because the proposition only precludes learning the “true” path-finding algorithm, not spurious memorizing solutions.
  3. 3.
    Small path length. If the path length $l$ is small, then either the search space of algorithms (which is about as large as $|\mathcal{C}|^{l}$) becomes tractable.
    1. (a)
      For example, if $l=3$, the first node is the only intermediate node between the start and goal token. Here, the model can easily learn the “right-to-left” solution that the desired node is the only node preceding the goal node.
Remark 2.
@@ Appendix D Other Remarks
(Mechanism implemented by teacherless model) The (hypothetical) solution that the teacherless model must implement is a fairly difficult one to implement — yet the model surprisingly learns to implement it. Recall that our hypothesis is that the teacherless model automatically learns to fit the targets in the reverse order, since the path from ${v}_{\tt{goal}}$ is unique. This is indeed what we find in Fig 4, where the accuracies of the later tokens become higher earlier. Note though that this is a fairly difficult computation to implement. First, when the model predicts $v_{i}$, it must require the identity of $v_{i+1}$. However, this identity is not fed as input to the model, in the absence of the teacher. Thus the model must have computed $v_{i+1}$ and crucially, stored that in one of its its internal representations. Then, by induction, when predicting the first node $v_{1}$, the model must know the identity of all the other nodes in the path. In other words, the model must have (a) computed and (b) stored the whole solution in its hidden representations before it outputs the first token. This is a substantial type of lookahead that some of our models are able to achieve under teacherless training.
Appendix E Experiment Notations
E.1 Verifying In-Distribution Performance
We simply compute exact match with the ground truth path as follows:
$\texttt{Acc}_{\texttt{ag}}(\texttt{LM}_{\theta}):=\mathbb{P}(\hat{\bm{r}}=\bm{r})\text{,}$ $\bm{p},\bm{r}\sim\mathcal{D},\hskip 5.69054pt\hat{\bm{r}}\stackrel{{\scriptstyle\mathclap{\mbox{{\tiny\tt ag}}}}}{{\sim}}\texttt{LM}_{\theta}.$ (5)
E.2 Quantifying the Clever Hans Cheat
$\hskip 5.69054pt\text{where}\hskip 8.53581pt\bm{p},\bm{r}\sim\mathcal{D},\hskip 5.69054pt\hat{\bm{r}}_{1<}\stackrel{{\scriptstyle\mathclap{\mbox{{\tiny\tt ag}}}}}{{\sim}}\texttt{LM}_{\theta}(\cdot;\bm{p},{v}_{\tt{start}},v_{1}^{\prime})$
E.3 Quantifying the Indecipherable Token Failure
To quantify the Indecipherable Token failure, in our path-finding task, we measure the accuracy in predicting the first token after the start node.
$\texttt{Acc}_{\texttt{1st}}(\texttt{LM}_{\theta})=\mathbb{P}\left(\hat{r}_{1}=r_{1}\right),$ $\bm{p},\bm{r}\sim\mathcal{D},\hat{\bm{r}}\stackrel{{\scriptstyle\mathclap{\mbox{{\tiny\tt ag}}}}}{{\sim}}\texttt{LM}_{\theta}(\cdot;\bm{p}).$ (7)
E.4 Inference in Teacherless Training
In teacherless training, we make use of an uninformative input $\bm{r}^{\$}$ that simply corresponds to a series of dummy tokens denotes by \$currency-dollar\\$\$. During inference, instead of autoregressing on the model’s own output, we use this uninformative input. We formalize this below:
$\hat{\bm{r}}\stackrel{{\scriptstyle\mathclap{\mbox{{\tiny\tt\$}}}}}{{\sim}}\texttt{LM}_{\theta}^{\$}(\cdot;\bm{p})\text{ where }\hat{r}_{i}\sim\texttt{LM}_{\theta}^{\$}(\cdot;\bm{p},\bm{r}^{\$}_{<i}).$ (8)
We then denote the accuracy of the model as follows:
$\texttt{Acc}_{\texttt{\$}}(\texttt{LM}_{\theta}^{\$})=\mathbb{P}\left(\hat{\bm{r}}=\bm{r}\right)$ $\bm{p},\bm{r}\sim\mathcal{D},\hskip 5.69054pt\hat{\bm{r}}\stackrel{{\scriptstyle\mathclap{\mbox{{\tiny\tt\$}}}}}{{\sim}}\texttt{LM}_{\theta}^{\$}(\cdot;\bm{p}).$ (9)
E.5 Reversed Training
Notationally, in reversed training we let $\texttt{LM}_{\theta}^{\texttt{rev}}$ be the model trained to maximize $\mathcal{J}_{\texttt{next-token}}$ with the targets (and the teacher-forced inputs) set to $\bm{r}^{\texttt{rev}}=r_{L_{\texttt{resp}}},\ldots r_{1}$, the reversal of $\bm{r}$. We then measure the autoregressive accuracy by comparing against $\bm{r}^{\texttt{rev}}$:
$\texttt{Acc}_{\texttt{rev}}(\texttt{LM}_{\theta}^{\texttt{rev}})=\mathbb{P}\left(\hat{\bm{r}}=\bm{r}^{\texttt{rev}}\right),$ $\;\bm{p},\bm{r}\sim\mathcal{D},\hat{\bm{r}}\stackrel{{\scriptstyle\mathclap{\mbox{{\tiny\tt ag}}}}}{{\sim}}\texttt{LM}_{\theta}^{\texttt{rev}}(\cdot;\bm{p})$ (10)
Appendix F More Experimental Results
F.1 Snowball Failure
@@
F.2 Clever Hans Cheating Accuracies
@@
F.3 More Detailed Accuracies
@@
F.4 Arithmetic Tasks
To further highlight that the identified failure modes are relevant for realistic tasks, we study the task of addition. We consider 3-digit addition, where samples are of the form
$x_{1}x_{2}x_{3}+y_{1}y_{2}y_{3}=z_{1}z_{2}z_{3}z_{4}$
We use the natural encoding (i.e. use the digits themselves as encodings) and pad shorter numbers with leading zeros to ensure fixed lengths. It has been observed in prior work Lee et al. (2023); Shen et al. (2023) that reversing the result (i.e. $z_{4}z_{3}z_{2}z_{1}$) is very beneficial for training, leading to significantly more sample-efficient learning. Here we study if our teacherless training strategy can lead to similar gains over the standard encoding, i.e. we encode the inputs as
$x_{1}x_{2}x_{3}+y_{1}y_{2}y_{3}=\$\$\$\$$
while keeping all other aspects of the training pipeline (such as the targets and the objective) the same as in standard training. We again train a GPT-Mini-style transformer with AdamW using a learning rate of $5$e-$4$. We display the resulting test accuracies for both standard, reversed and teacherless training in Fig. 6, as a function of training iterations. Here we follow previous works and at every iteration, sample with replacement from all the possible $3$-digit configurations (aside from a test set put aside before). We can indeed see that teacherless training leads to more sample-efficient learning compared to the standard format, but to slightly less efficient learning compared to the reverse format. This again hints at the fact that some form of CleverHans cheating is picked up when learning addition.
In this case though, it is more difficult to precisely characterize the form of the cheat in contrast to the path-finding problem. One possible form of cheating may be that in addition, knowing the leading digits of the answer can provide information about the subsequent digits. For example, when we add two single-digit numbers (picked uniformly at random), knowing that the $10^{\prime}th$ place is a $1$ rather than a 00, tells us that the 00th place is more likely to be a smaller digit like 00.
Figure 6: Test accuracies for 3-digit addition for standard, reversed and teacherless training.
F.5 Predicting Only the Indecipherable Token
To study the difficulty of the first token in isolation from the Clever Hans cheat, we reduce the task to solely predicting this token, removing the rest of the path. The problem thus effectively becomes a classification problem. We use the same learning setup as for previous experiments and keep the same number of examples. We display the results in Table 5. Our results indicate that the model fails to learn the first token in isolation.
This reinforces our understanding of the role played by the Clever Hans cheat in bringing about failure. One possibility could have been that the first node in isolation is easy to learn, but the Clever Hans cheat makes it harder somehow (e.g., it attracts the model to a bad local minimum). The other possibility (the one we put forth) is that the Clever Hans cheat reduces the problem to learning the first token in isolation, which we argued is hard. These experiments confirm the second possibility.
$G_{2,5}$ $G_{2,20}$ $G_{5,5}$ $G_{10,5}$ $G_{20,5}$
GPT2-Large $50.2$ $50.4$ $18.9$ $10.4$ $4.5$
Table 5: Accuracy when solely learning to predict the difficult token $r_{1}$
Appendix G Other experimental details
G.1 Tokenization
@@ G.2 Models
Appendix H More Related Work
Other arguments about next-token prediction. We note that the works of Kääriäinen (2006); Ross & Bagnell (2010) capture a stronger notion of snowballing, wherein, once an erroneuous sub-optimal action is committed, the model is more likely to commit more sub-optimal actions since it has wandered into territories that it was not trained on. Implicitly, the error here is not evaluated as an exact match of the response (i.e., $\bm{r}\neq\hat{\bm{r}}$) but as a cumulative notion of error over all steps (e.g., $\sum\mathbf{1}[r_{i}\neq\hat{r}_{i}]$). In this setting, there is an additional cause of failure called exposure bias: the teacher-forced model has only been trained on correct trajectories, and has not learned how to recover from poor trajectories. Nevertheless, even this notion of snowballing assumes that that teacher-forcing has learned an accurate next-token predictor in the first place, which our failure mode challenges.
A closely-related criticism (Bubeck et al., 2023; Dawid & LeCun, 2023; LeCun, 2024; Du et al., 2023a) is that to model human thinking, we need to model two types of thinking as outlined in Kahneman (2011): a fast (System 1) thinking process that is also guided by a slower (System 2) thinking process. Theoretically, Lin et al. (2021) show that there are formal languages for which expressing some next-tokens may require super-polynomial time or parameter count during inference. These arguments however only suggest that some tokens require more computation; not that they are specifically problematic under left-to-right learning. However, Du et al. (2023a) informally note that some next tokens can be hard to learn as they require a global understanding of what will be uttered in the future (but see Remark 3 below).
Remark 3.
(Locally Unlearnable Token vs Indecipherable Token) We note that the “locally unlearnable” hypothesis of Du et al. (2023a) is related to, but not the same as the Indecipherable Token failure. The hypothesis in Du et al. (2023a) is that when we learn tokens left-to-right, some tokens simply cannot be learned since crucial information becomes available only in subsequent tokens. This hypothesized failure may happen regardless of whether the supervision from subsequent tokens is lost to a Clever Hans cheat. In contrast, in our path-star graph, the Indecipherable Token becomes unlearnable only because of the Clever Hans cheat. For example, the (first) Indecipherable Token in the path-star problem is locally learnable by the teacherless model (where, the crucial information is still only presented after this token). This token becomes unlearnable only in the teacher-forced model where the Clever Hans cheat emerges.
@@ Appendix H More Related Work
We survey related arguments of next-token prediction, orthogonal to our main discussion regarding planning. Allen-Zhu & Li (2023); Lv et al. (2023) report that language models that are trained on A equals B are unable to infer B equals A, which Allen-Zhu & Li (2023) suggest is due to autoregressive left-right training. Du et al. (2023b); Welleck et al. (2020) formalize the limitation that autoregressive models may potentially assign non-zero probability to infinite-length strings, thus leading to non-terminating inference. Li et al. (2024) provide a Transformer-specific analysis of how self-attention affects the optimization geometry of next-token prediction. Thrampoulidis (2024) provide an analysis of the implict bias of optimization with next-token prediction for linear models.
@@
Other limitations of Transformers Merrill & Sabharwal (2023) identify limitations of the representative power of Transformer architecture when the arithmetic precision is logarithmic in the number of input tokens. Bender et al. (2021) criticize GPT-like language models as simply parroting out training data with minor stochasticity, while Arkoudas (2023) report that such models struggle with reasoning, even if not a stochastic parrot. Young & You (2022) study masked language (T5, BERT) models (not causally-trained) and argue there are inconsistencies in the probabilities that they assign. E.g., when conditioned on ‘white’, the probability of ‘rice’ may be higher ‘bread’ but the probability of ‘white bread’ and ‘white rice’ are the opposite. Artetxe et al. (2022) empirically analyze the effect of bidirectional attention and bidirectional supervision (as in masked language modeling) during pretraining on the ability of the model to do various things, including next-token prediction. Springer et al. (2024) argue that autoregressive Transformers compute sub-optimal embeddings that can be improved by repeating the input text twice.
@@
Such limits of end-to-end learning have been echoed in literature on learning with chain-of-thought-type supervision. Recent theoretical works have shown broad classes of tasks (e.g., any function efficiently computed by a Turing machine) where prepending CoT to the end target allows efficiently learning tasks; yet, there are “multi-hop reasoning” tasks that are unlearnable end-to-end (i.e., without intermediate supervision) either due to computational hardness (Wies et al., 2023) or representational limits (Malach, 2023)). Earlier theoretical works Shalev-Shwartz et al. (2017); Shalev-Shwartz & Shashua (2016) have similarly proven negative results for end-to-end learning in similar settings. Similar empirical arguments have been made in neural network literature (Gülçehre & Bengio, 2016; Glasmachers, 2017) and also more recently, in language models on complex reasoning and math problems (Nye et al., 2021; Ling et al., 2017; Cobbe et al., 2021; Piekos et al., 2021; Zelikman et al., 2022; Recchia, 2021; Cobbe et al., 2021; Hsieh et al., 2023; Shridhar et al., 2022).
Remark 4.
(Chain-of-thought before vs. after end target.) It is worth noting though that the above lines of work are concerned with chain-of-thought that is present before the end target; in our setup, this supervision is presented only after the end target. Surprisingly, some of our teacherless models manage to utilize even such hindsight chain-of-thought. This success is not fully explained by existing positive results about chain-of-thought supervision, such as Wies et al. (2023); Malach (2023), where supervision is provided before the end target.
Going beyond next-token prediction. Inference-time techniques like chain-of-thought (Reynolds & McDonell, 2021; Wei et al., 2022; Kojima et al., 2022) and its variants (Yao et al., 2023a; Besta et al., 2024; Yao et al., 2023b) or those that elicit feedback from the model (Madaan et al., 2024; Huang et al., 2022; Shinn et al., 2023) can be thought of as going beyond conventional form of inference by allowing the model to think more before producing its final answer. However, the backbone in these models are still trained by standard teacher-forcing. While other techniques (Burtsev et al., 2020; Xue et al., 2023; Goyal et al., 20234) train the model to explicitly think more, even these boil down to next-token prediction during training.
One may argue that reinforcement learning-based training (Ranzato et al., 2016; Wu et al., 2016; Bahdanau et al., 2017; Paulus et al., 2018; Ziegler et al., 2019; Liu et al., 2020; Ouyang et al., 2022) is another way to build backbones that go beyond teacher-forcing. However, it is worth noting that the gradients in these techniques boil down to teacher-forcing on the model’s own generated answer. Furthermore, if we desire that the model be able to generate a solution that can plan ahead of time, it is unclear how a model can go from a complete inability to plan (that may assign near-zero probability to the true plan in an exponential space of solutions), to discovering the correct plan simply through preference-based feedback (see (Havrilla et al., 2024) for related empirical evidence).
Another line of work — spanning language (Bengio et al., 2015; Goyal et al., 2016), imitation learning (Ross et al., 2011; Ross & Bagnell, 2010, 2014) and structured prediction (Daumé III et al., 2009; Chang et al., 2015) — has been aimed at addressing the Snowball Failure, under the assumption that the model has otherwise learned an accurate next-step predictor. Broadly, the idea is to train the model on a mixture of the ground truth sequences and the model-generated sequences themselves, as a way to ensure that the test-time and training-time distributions are as similar as possible. These techniques however do not address the failure to learn a good next-step predictor in the first place.
@@
Predicting future tokens. Some works (Gurnee et al., 2023; Meng et al., 2022; Pal et al., 2023) aim to recover future tokens that an already-trained model may predict based on the internal layers of the current token. Note that the success of this does not imply that the model necessarily plans well. This only means that it is possible to recover what the already-trained model wants to generate in the future (which may simply be a bad plan). Pfau et al. (2023) train a language model to predict in reverse with the orthogonal goal of finding prefixes that elicit certain behaviors.
Shortcut-learning in language models. A line of work has empirically and theoretically analyzed how Transformer-based language models learn superficial shortcuts to (partially) solve tasks such as learning multiplication (Dziri et al., 2024), logic (Zhang et al., 2023), automata (Liu et al., 2023), recursion (Young & You, 2022), reading comprehension (Lai et al., 2021) and multiple-choice questions (Ranaldi & Zanzotto, 2023) However, these shortcuts must not be confused with the Clever Hans cheating induced by teacher-forcing as elaborated below.
Remark 5.
(Difference between Clever Hans cheating and known shortcut-learning failures in Transformers.) First, these aforementioned shortcuts exist independent of teacher-forcing: these are correlations between the prefix (such as the initial digits of two multiplicands) and the final answer (the initial digits of the product) in the underlying training distribution. But Clever Hans cheats arise only upon teacher-forcing: these are correlations between the prefixes of the answer itself to the rest of the answer. Second, the above shortcuts only fail out-of-distribution (such as when the number of multiplied digits is increased, where the failure is in length generalization (Anil et al., 2022)). In contrast, the Clever Hans cheat is more severe as it causes in-distribution failure. Thirdly, the aforementioned empirical observations are specific to Transformers, and the theoretical arguments rely crucially on properties of the Transformer (such as its non-recurrence and convolution, or its self-attention modules). Our argument however only relies on the teacher-forcing objective with no reliance on the Transformer architecture, and is demonstrated even for the recurrent Mamba architecture.